RECENT APPLICATIONS AND FUTURE PROSPECTS OF METHANE SPECTROSCOPY TO THE AT-MOSPHERE OF TITAN

\underline{V. BOUDON}, Institut Carnot de Bourgogne, UMR 5209 CNRS-Université de Bourgogne, 9. Av. A. Savary, BP 47870, F-21078 Dijon Cedex, France; A. COUSTENIS, E. LELLOUCH, P. DROSSART, LESIA, Observatoire de Paris-Meudon, 5 Place Jules Jansen, F-92195 Meudon Cedex, France; A. NEGRÃO, Observatorio Astronomico and Faculdade de Ciências da Universidade de Lisboa, 1749-016 Lisboa, Portugal; C. A. GRIFFITH, Lunar and Planetary Laboratory, University of Arizona, Tucson, AZ, 85721 USA.

The advent of the Cassini-Huygens mission studying Saturn’s system and its largest moon Titan, has led to a renewed interest in reliable models of the absorption spectrum of methane. CH_{4} is the main absorber in Titan’s thick atmosphere. Although the models developed in the Dijon group still do not allow sufficently reliable simulations above 5000 cm^{−1} to reproduce all the recent data (such as Huygens/DISR spectra, for instance), the methane coefficients in the 0–4800 cm^{−1} region have contributed to a better understanding of various ground- and spaced-based data : ISO high-resolution data in the 3 μm region^{a}, near-infrared VLT data^{b} and 2 μm VLT data of the Huygens probe landing site^{c}. These coefficients have also contributed to the discovery of a polar ethane cloud on Titan^{d}. Models of CH_{4} hot bands by the Dijon group are also of primary importance for fluorescence calculations observed at 3.3 μm with Cassini/VIMS^{e}. After a short review of these works, future prospects for line-by-line analyses of CH_{4} spectra for planetary applications will also be discussed and compared to other approaches, such as the so-called band models or purely experimental approaches.

aA. Coustenis, A. Negrão, A. Salama et al. Icarus 180, 176–185 (2006).

bA. Negrão, A. Coustenis, E. Lellouch et al. Planet. Space Sci. 54, 1225–1246 (2006).

cA. Negrão, M. Hirtzig, A. Coustenis et al. J. Geophys. Res. Planets 112, E02S92 (2007).

dC. A. Griffith, P. Penteado, P. Rannou et al. Science 313, 1620–1622 (2006).

eR. H. Brown, K. H. Baines, G. Bellucci et al., Astron. Astrophys. 446, 707–716 (2006).